The cylinder of a pump has a volume of 25.0 cm3. It is connected to a container of


volume 225 cm3, containing air at atmospheric pressure (100 kPa). The pump draws air


from the atmospheric and pumps it into the container. Assume that the pump stroke is


slow and therefore the temperature of the air in the container remains constant.


Calculate the pressure of the air in the flask after one stroke of the pump adds more air.

1
Expert's answer
2021-12-29T12:35:03-0500

Apply the universal gas law keeping in mind the process did not affect the temperature and volume:


"PV=\dfrac mMRT,\\\space\\
P_1V=\dfrac {m_1}MRT."


Divide one by another:


"\dfrac{P}{P_1}=\dfrac{m}{m_1}=\dfrac{m}{m+\Delta m},\\\space\\
P_1=P\dfrac{m+\Delta m}{m}=P\dfrac{v+\Delta v}{v}=111\text{ kPa}."

*We used "m=v\rho." Since the density of gas was constant while it was taken from the atmosphere, the density is equal everywhere.